OpenStudy (republican31):

I need a formula to solve for x. y = x/3 + 8

OpenStudy (anonymous):

1/3x+8=y 8 doesn't contain the variable so move it to the right subtract 8 from both sides 1/3x=-8+y 1/3x=y-8 multiply by 3 x=y*(3)-8*(3) x=y(3)-8(3) x=3y-8(3) multiply -8 x=3y-24

OpenStudy (anonymous):

thanks